OpenStudy (anonymous):

In how many ways can 10 instructors be assigned to ten sections of a course in mathematics?

OpenStudy (anonymous):

10!

OpenStudy (anonymous):

\[10\times 9\times 8\times 7\times 6\times 5\times 4\times 3\times 2\times 1\]